# Look into the Future: Deep Contextualized Sequential Recommendation

### v1

@@ Abstract.
Sequential recommendation focuses on mining useful patterns from the user behavior history to better estimate his preference on the candidate items. Previous solutions adopt recurrent networks or retrieval methods to obtain the user’s profile representation so as to perform the preference estimation. In this paper, we propose a novel framework of sequential recommendation called Look into the Future (LIFT), which builds and leverages the contexts of sequential recommendation. The context in LIFT refers to a user’s current profile that can be represented based on both past and future behaviors. As such, the learned context will be more effective in predicting the user’s behaviors in sequential recommendation. Apparently, it is impossible to use real future information to predict the current behavior, we thus propose a novel retrieval-based framework to use the most similar interaction’s future information as the future context of the target interaction without data leakage. Furthermore, in order to exploit the intrinsic information embedded within the context itself, we introduce an innovative pretraining methodology incorporating behavior masking. This approach is designed to facilitate the efficient acquisition of context representations. We demonstrate that finding relevant contexts from the global user pool via retrieval methods will greatly improve preference estimation performance. In our extensive experiments over real-world datasets, LIFT demonstrates significant performance improvement on click-through rate prediction tasks in sequential recommendation over strong baselines. The experiment code is provided at https://anonymous.4open.science/r/LIFT-277C/.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06conference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management ; October 21 –– 25 2024 ; Barcelona, Spain
@@ 1. Introduction
As illustrated in Figure 1, the major techniques for deep learning-based sequential recommendation are generally threefold. Firstly, deep architectures are designed to extend the capability of traditional collaborative filtering models for better mining of the feature interactions (Guo et al., 2017; Zhang et al., 2016; Wang et al., 2017). Secondly, different models like recurrent neural networks (RNN), memory networks, or transformers are adopted to learn an effective representation of the user behavior sequence, based on which a deep learning predictor is built to predict the behavior label for the current-step recommendation (Pi et al., 2019; Chen et al., 2018; Sun et al., 2019). Dealing with the long behavior sequence problem, thirdly, the retrieval methods are leveraged to fetch far early but relevant behaviors, which are then aggregated according to the target prediction condition and fed into the final label predictor (Qi et al., 2020; Qin et al., 2020).
However, the retrieval-enhanced models for sequential recommendation still retrieve in the “raw” data space, i.e., the individual behavior data, which actually loses the contextual information of the user making that particular behavior. Users’ behaviors over the recommended items commonly exhibit interdependence. Typically, a user browses items with a specific intent, and within a certain time window, his attention is often drawn to items within a particular theme. For example, when a user is browsing ketchup, they may also browse many other kitchen condiments before and after. To this end, we posit the concept of a user’s browsing sequence encompassing historical and future behaviors as the context of that behavior. The incorporation of contextual information pertaining to the future behavior enhances the effectiveness of modeling user behaviors. Similarly, in the area of natural language processing (NLP), to model the word representation, the words before and after the target word are often used (Mikolov et al., 2013; Devlin et al., 2018). Instead, in the field of recommendation systems, the future behavior sequences are often overlooked (Zhou et al., 2018b, 2019) because the future information is inaccessible when predicting the current behavior. However, we may obtain insight into future user behavior by identifying analogous contexts from the observed data of other users and employing them to better model the context of the target user.
@@
In this paper, we propose a novel framework of sequential recommendation called Look into the Future (LIFT), which focuses on retrieval in the user contextual information space while leveraging the future behaviors of the retrieved context, as illustrated on the bottom part of Figure 1. LIFT works in scenarios where the behavior sequences of multiple users are accessible. In sequential recommendation, the representation of context interactions plays an important role in behavior prediction. All previous works in recommendation focus on modeling the historical part of the context. From a hindsight perspective, the entire context of the target interaction should include future interactions. As mentioned above, it is impossible to use the real future information to predict the current behavior. Thus, given the current-step context with the candidate item for behavior label prediction, LIFT performs retrieval to fetch the most similar interaction behavior from the whole user pool. Then, extending from each of the retrieved behaviors, the historical and future behavior data are imported to enrich the context of that retrieved behavior. Note that each of the included ‘future’ behaviors is still temporally earlier than the current timestep of the target user behavior, which avoids any data leakage issue. Such ‘future’ behaviors can be regarded as a kind of privileged information (Vapnik et al., 2015; Xu et al., 2020), which is not accessible for current-step prediction but is accessible for historical behavior data. To our knowledge, there is no previous work on sequential recommendation that performs effective retrieval of multiple users’ behavior contexts.
@@
Figure 1. The comparison between LIFT and conventional models entails several key distinctions: a) Traditional models rely solely on instant user and item information when making predictions. b) Sequential models, conversely, typically incorporate the user’s historical interactions to capture their evolving interests over time. c) Retrieval-based models perform retrieval to fetch far-before but relevant historical behaviors to build the user profile for predictions. d) LIFT focuses on interaction context, encompassing both the historical and future sequence of interactions for each user-item interaction. LIFT’s primary objective is to exploit this comprehensive context information to enhance prediction accuracy. However, note that in practical scenarios, accessing genuine future information is unattainable. To overcome this limitation, we leverage the future sequences of relevant observed interactions as a surrogate to approximate the future context for the current interaction. Note that the included ‘future’ interactions are still temporally earlier than the current target behavior for prediction, which avoids any data leakage issue.
@@
Furthermore, it is crucial not only to utilize information derived from the target sample but also to exploit the inherent information within the context itself for the purpose of enhancing contextual representation. Therefore, besides the supervised training with the label prediction loss, a representation learning loss is much important in our task. To learn an effective representation of the context with both history and future behavior data, we further devise a pretraining method that performs masked behavior prediction during the pretraining stage. We conduct extensive experiments on three real-world sequential recommendation datasets, where LIFT demonstrates significant performance improvements over strong baselines.
@@
  • We propose to incorporate the future information as part of the context information. We propose a novel framework that leverages retrieval techniques to use relevant context information to enhance the prediction performance of sequential recommendations while still avoiding data leakage. To our knowledge, this is the first work that makes the most use of the entire context information in the recommendation system based on retrieval techniques.
@@
  • To prevent noise introduced by the retrieved data, we introduce a key-based attention aggregation technique aimed at consolidating contextual representations in accordance with the specific target sample and the keys contained within the retrieved contextual information.
@@ 2. Related Work
Sequential Recommendation. For sequential recommendation, user behavior modeling is the core technique that mines user preferences from historical interaction behaviors meticulously. To better extract informative knowledge from user’s behavior sequence, various network structures have been proposed, including Recurrent Neural Networks (RNNs), Convolutional Neural Networks (CNNs), Attention Networks, and Memory Networks. GRU4Rec (Hidasi et al., 2015) designs Gated Recurrent Units (GRUs) to capture the preference-evolving relationship while Caser (Tang and Wang, 2018) leverages the horizontal and vertical convolution to model skip behaviors at both union-level and point-level. Moreover, the attention mechanism is the most popular method for modeling item dependencies, and several influential works are proposed, including SASRec (Kang and McAuley, 2018), DIN (Zhou et al., 2018b), DIEN (Zhou et al., 2019), and BERT4Rec (Sun et al., 2019). Among these, DIN and DIEN leverage the target-attention network to identify important historical items, while SASRec and BERT4Rec apply the self-attention network with Transformer architecture to excavate behavior dependencies. Besides, memory-based methods (Wang et al., 2018; Pi et al., 2019) are also proposed to store user behavior representations in a read-write manner. To better mine user preferences from their historical behaviors, two kinds of work are proposed to further extend the behavior modeling. The first one is multi-behavior modeling (Zhou et al., 2018a; Yuan et al., 2022) that explicitly leverages different types of behavior (e.g., click and purchase behaviors) to measure item correlations within different behavior sequences, thus capturing users’ diverse interests. The second one is behavior modeling with side information (Zhang et al., 2019) that involves various item attribute features (e.g., category) in addition to item IDs for better exploiting the rich knowledge.
@@
Pretraining for Recommendation. Pretraining the deep learning models (or the data representation) with self-supervised learning methods has been widely studied in natural language processing and computer vision. Generally, there are two major categories of self-supervised training methods, namely contrastive learning (Gao et al., 2021; Chen et al., 2020) and mask recovery (Brown et al., 2020; Jacob et al., 2019; He et al., 2022). For recommender systems or tabular data prediction tasks, there are some recent attempts in this direction. To list a few examples, BERT4Rec (Sun et al., 2019) focuses on learning the representation of sequential behaviors via masking the final item of a subsequence of user behaviors. SCARF (Bahri et al., 2022) raises a contrastive learning loss via feature corruption over the tabular data. MISS (Guo et al., 2022) proposes interest-level contrastive losses to take the place of sample-level losses in order to mine self-supervision signals from user behaviors of multiple interests. Also, there are recent attempts to leverage the knowledge from the pretraining work on outsourced data, such as knowledge graphs (Wong et al., 2021) and language corpus (Cui et al., 2022), to enhance the recommender systems.
@@ 3. Formulation & Preliminaries
Notation Description
$x_{z}$, $\mathbf{x}_{z}$ The raw feature and the embedding of $z$-th sample
$x_{t}$, $\mathbf{x}_{t}$ The raw feature and the embedding of target sample
$E_{\omega}$ The encoder and its parameters
$R$ The retriever
$F_{\theta}$ The predictor and its parameters
$d$ An interaction
$v$ The encoder output vector dimension
$w$ The feature embedding vector dimension
$K$ The size of retrieved contexts
$L$ The length of history and future sequence
$c$, $h$, $f$ The context sequence, history sequence, future sequence
Table 1. Notations and corresponding descriptions.
In this section, we formulate the studied problem and introduce the preliminaries. The notations and their descriptions are summarized in table 1. A recommendation dataset $\mathcal{D}=\{d_{z}\}_{z=1}^{N}$ is represented as a set of interactions $d_{z}$ between user $u_{z}$ and item $i_{z}$. Each interaction $d_{z}$ can be formulated as the feature-label pair $(x_{z},y_{z})$, where $x_{z}=\{x_{j}^{z}\}_{j=1}^{M}$ contains the user and item features and a label $y\in\{0,1\}$ indicating whether the user will click the item or not, i.e., click-through rate (CTR).
@@ 4. The LIFT Framework
Figure 3. The overview of major components and workflow in LIFT. In the initial phase (Stage 1), we extract a sequence dataset from the original dataset, subsequently employing this sequence dataset for the purpose of pretraining the encoder. In the subsequent phase (Stage 2), we establish a datastore utilizing the dataset $\mathcal{D}_{\text{retrieval}}$. Following this, we proceed to train the predictor denoted as ”$Predictor$” using the dataset $\mathcal{D}_{\text{train}}$.
@@ 4.1. Overview
There are two steps in LIFT framework: pretraining and predicting. During pretraining, an encoder $E$ is trained on a sampled $\mathcal{S}_{\text{pretrain}}$ from $\mathcal{D}_{\text{retrieval}}$ For $predicting$, the LIFT model is first using the retriever $R$ to find the top-$K$ relevant user context sequences $C_{t}$. And then, each sequence in $C_{t}$ will be encoded by $E_{\omega}$. After aggregating the information of the target user and item $x_{t}$, the target user’s history $h_{t}$, retrieving sequence $C_{t}$, the predictor $F_{\theta}$ will generate the output label prediction.
@@ 4.2. Encoder
Model Architecture. LIFT adopts a decoder-only Transformer as the encoder, which is based on the original implementation described in Vaswani et al. (2017). Transformer has been shown to perform strongly in language modeling (Brown et al., 2020; Devlin et al., 2019), machine translation (Vaswani et al., 2017), recommender system (Sun et al., 2019), etc. Different tasks may leverage different parts of the Transformer architecture. The difference between the Transformer encoder and decoder is that the self-attention sub-layer in the decoder is modified to make sure that the prediction in position $p$ could only depend on the positions earlier than $p$ (Vaswani et al., 2017). For the data of the recommender system exhibiting temporal characteristics, we use the decoder part of the Transformer. We also compare the performance of different encoders in the experiment, which shows the superiority of the Transformer decoder.
@@
To feed the interaction $d_{z}$ into the encoder, $d_{z}$’s feature part $x_{z}$ is fed into an embedding layer. The embedding layer will map the one-hot encoded vector $x_{z}$ into a dense vector $\bf{x}_{z}$, where each feature $a_{i}$ in $x_{z}$ will be mapped into a $w$-dimensional vector and then be concatenated into one $(M\times w)$-dimensional dense vector. The label part of the $d_{z}$ will be preserved. In the training stage, if the interaction $d_{z}$ is chosen to be masked, the label will be replaced by a special token [MSK]. We will describe the training loss in the next subsection. After processing the sequence, the output will be sent into the decoder-only Transformer decoder sequentially, and the output is a set of hidden states $O=\{o_{1},o_{2},...,o_{T}\}$. We choose the last state $o_{T}$ as the representation of the interaction sequence $\bf{s}$.
Data. The pretraining dataset $\mathcal{S}_{\text{pretrain}}$ is sampled from $\mathcal{D}_{\text{retrieval}}$. For each user $u$ in $\mathcal{U}$, we use all the interactions in $\mathcal{D}_{retrievel}$ of $u$ to assemble an interaction sequence $s_{u}$. Then we get a sequence set $S$. For any sequence $s$ in $S$, we pick up a subsequence every $L$ interactions. Using these subsequences, we get the pretraining dataset $\mathcal{S}_{\text{pretrain}}$.
Mask Behavior Loss. To leverage the inherent information contained within the context, we formulate a self-supervised loss function. In contrast to certain existing sequential recommendation models (Zhou et al., 2018b, 2019), wherein only the target interaction label is considered as the prediction target for any given interaction sequence, all other interaction labels within the sequence are omitted. It is rational to argue that incorporating more behavioral information from the context as the supervision signal would lead to a more comprehensive representation of the context. Consequently, such an approach is expected to yield greater benefits for the prediction of target behavior.
As shown in Figure 2, the left part describes the encoder architecture and loss in the pretraining procedure. In order to train a deep representation of the sequence, we mask part of the interaction’s labels and then predict those masked behavior. Unlike the Cloze task (Taylor, 1953) in natural language processing, it is difficult to predict the masked tokens in the user interaction sequences because there is no syntax in the recommendation. An interaction $d_{i}$ in the input sequence $s$ can be represent as
@@
We adopt the binary cross entropy loss as the objective function. In the pretraining procedure, we feed the masked interactions’ corresponding output as input into a multi-layer perceptron (MLP) to predict the masked label. Thus the pretrain loss $\mathcal{L}_{\text{pretrain}}$ can be represented as $\mathcal{L}_{\text{pretrain}}=$
(10) $-\sum_{d_{i}\in\mathcal{M}}\left[y_{i}\log\sigma(\text{MLP}(o_{i}))+(1-y_{i})\log(1-\sigma(\text{MLP}(o_{i})))\right],$

### v2

@@ Abstract.
Sequential recommendation aims to estimate how a user’s interests evolve over time via uncovering valuable patterns from user behavior history. Many previous sequential models have solely relied on users’ historical information to model the evolution of their interests, neglecting the crucial role that future information plays in accurately capturing these dynamics. However, effectively incorporating future information in sequential modeling is non-trivial since it is impossible to make the current-step prediction for any target user by leveraging his future data. In this paper, we propose a novel framework of sequential recommendation called Look into the Future (LIFT), which builds and leverages the contexts of sequential recommendation. In LIFT, the context of a target user’s interaction is represented based on i) his own past behaviors and ii) the past and future behaviors of the retrieved similar interactions from other users. As such, the learned context will be more informative and effective in predicting the target user’s behaviors in sequential recommendation without temporal data leakage. Furthermore, in order to exploit the intrinsic information embedded within the context itself, we introduce an innovative pretraining methodology incorporating behavior masking. In our extensive experiments on five real-world datasets, LIFT achieves significant performance improvement on click-through rate prediction and rating prediction tasks in sequential recommendation over strong baselines, demonstrating that retrieving and leveraging relevant contexts from the global user pool greatly benefits sequential recommendation. The experiment code is provided at https://anonymous.4open.science/r/LIFT-277C/Readme.md.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06conference: Proceedings of the 31th International ACM SIGKDD Conference on Knowledge Discovery & Data Mining; August 3–7, 2025; Toronto, ON, Canada
@@ 1. Introduction
As illustrated in Figure 1, the major techniques for deep learning-based sequential recommendation are generally threefold. Firstly, deep architectures are designed to extend the capability of traditional collaborative filtering models for better mining of the feature interactions (Guo et al., 2017; Zhang et al., 2016; Wang et al., 2017). Secondly, different models like recurrent neural networks (RNN), memory networks, or transformers are adopted to learn an effective representation of the user behavior sequence, based on which a deep learning predictor is built to predict the behavior label for the current-step recommendation (Pi et al., 2019; Chen et al., 2018; Sun et al., 2019). Dealing with the long behavior sequence problem, thirdly, the retrieval methods are leveraged to fetch far early yet relevant behaviors, which are then aggregated according to the target prediction condition and fed into the final label predictor (Qi et al., 2020; Qin et al., 2020).
The primary aim of sequence modeling in recommender systems is to capture the evolving trends in user interests (Zhou et al., 2019). Merely relying on historical data is insufficient to model these dynamic changes effectively (Yuan et al., 2020). Incorporating future information provides a hindsight perspective to help predict shifts in user preferences. For instance, after purchasing a smartphone, a user is likely to show significant interest in related accessories. However, incorporating future data for predictive modeling without causing data leakage presents a significant challenge. Most existing models (Zhou et al., 2018b; Kang and McAuley, 2018) neglect the potential of future user behaviors, while a few attempts (Yuan et al., 2020; Sun et al., 2019) implicitly use future information during training but fail to leverage it during the inference stage, leaving a significant potential for improvement.
To address this issue, we employ retrieval-based methods to introduce context that contains future information into the model. Previously, retrieval techniques have been integrated into sequential recommendations to effectively access and leverage longer historical information (Pi et al., 2019; Qin et al., 2020). Nonetheless, these retrieval methods typically focus on raw data from history, overlooking the importance of context. The interaction sequences of similar users with similar items tend to follow comparable patterns. By utilizing retrieval methods, we can fetch similar contexts in the log data from other users to approximate the future information of the target user. This approach eliminates the issue of temporal data leakage while enhancing the model’s ability to predict user behaviors.
@@
In this paper, we propose a novel framework of sequential recommendation called Look into the Future (LIFT), which focuses on retrieval in the user contextual information space while leveraging the future behaviors of the retrieved contexts, as illustrated on the bottom part of Figure 1. As mentioned above, it is infeasible to use real future information to predict the current behavior. Thus, given the current-step context with the candidate item for behavior label prediction, LIFT performs retrieval to fetch the most similar interaction behavior from the whole user pool. Then, extending from each of the retrieved behaviors, the historical and future behavior data are imported to enrich the context of that retrieved behavior. Note that each of the included future behaviors is still earlier than the current timestep of the target user behavior, which avoids any data leakage issue. Such future behaviors can be regarded as a kind of privileged information (Vapnik et al., 2015; Xu et al., 2020), which is not accessible for current-step prediction but is accessible for historical behavior data. To our knowledge, there is no previous work on sequential recommendation that performs effective retrieval of multiple users’ behavior contexts that include both past and future behaviors.
@@
Figure 1. The comparison between LIFT and conventional models entails several key distinctions: a) Traditional models rely solely on instant user and item information when making predictions. b) Sequential models, conversely, typically incorporate the user’s historical interactions to capture their evolving interests over time. c) Retrieval-based models perform retrieval to fetch far-before but relevant historical behaviors to build the user profile for predictions. d) LIFT focuses on interaction context, encompassing both the historical and future sequence of interactions for each user-item interaction.
@@
Furthermore, it is crucial not only to utilize information derived from the target sample but also to exploit the inherent information within the context itself for the purpose of enhancing contextual representation. Therefore, besides the supervised training with the label prediction loss, a representation learning loss is much important in our task. To learn an effective representation of the context with both history and future behavior data, we further devise a pretraining method that performs masked behavior prediction during the pretraining stage. Moreover, to reduce the noise introduced by the retrieved data, we designed an attention mechanism that assigns different weights to the retrieved data using dense embeddings of users and items. We conduct extensive experiments on five real-world sequential recommendation datasets with click-through rate prediction and top-N ranking tasks, where LIFT demonstrates significant performance improvements over strong baselines.
@@
  • We propose a novel LIFT framework that incorporates future information as part of the context. By leveraging retrieval techniques, LIFT utilizes relevant contextual information to enhance the prediction performance of sequential recommendations while avoiding temporal data leakage. To our knowledge, this is the first work to integrate future information comprehensively in both training and inference phases within a recommender system based on retrieval techniques.
@@
  • To avoid noise in the retrieved behavior sequences, we propose a key-based attention mechanism to aggregate the retrieved data.
@@ 2. Related Work
Sequential Recommendation. For sequential recommendation, user behavior modeling is the core technique that mines user preferences from historical interaction behaviors meticulously. To better extract informative knowledge from user’s behavior sequence, various network structures have been proposed, including Recurrent Neural Networks (RNNs), Convolutional Neural Networks (CNNs), Attention Networks, and Memory Networks. GRU4Rec (Hidasi et al., 2015) designs Gated Recurrent Units (GRUs) to capture the preference-evolving relationship while Caser (Tang and Wang, 2018) leverages the horizontal and vertical convolution to model skip behaviors at both union-level and point-level. Moreover, the attention mechanism is the most popular method for modeling item dependencies, and several influential works are proposed, including SASRec (Kang and McAuley, 2018), DIN (Zhou et al., 2018b), DIEN (Zhou et al., 2019), and BERT4Rec (Sun et al., 2019). To better mine user preferences from their historical behaviors, two kinds of work are proposed to further extend the behavior modeling. The first one is multi-behavior modeling (Zhou et al., 2018a; Yuan et al., 2022) that explicitly leverages different types of behavior (e.g., click and purchase behaviors) to measure item correlations within different behavior sequences, thus capturing users’ diverse interests. The second one is behavior modeling with side information (Zhang et al., 2019) that involves various item attribute features (e.g., category) in addition to item IDs for better exploiting the rich knowledge.
@@
Pretraining for Recommendation. Pretraining the deep learning models (or the data representation) with self-supervised learning methods has been widely studied in natural language processing and computer vision. Generally, there are two major categories of self-supervised training methods, namely contrastive learning (Gao et al., 2021; Chen et al., 2020) and mask recovery (Brown et al., 2020; Jacob et al., 2019; He et al., 2022). For recommender systems or tabular data prediction tasks, there are some recent attempts in this direction. To list a few examples, BERT4Rec (Sun et al., 2019) focuses on learning the representation of sequential behaviors via masking the final item of a subsequence of user behaviors. SCARF (Bahri et al., 2022) raises a contrastive learning loss via feature corruption over the tabular data. MISS (Guo et al., 2022) proposes interest-level contrastive losses to take the place of sample-level losses in order to mine self-supervision signals from user behaviors of multiple interests. S3(Zhou et al., 2020) proporsed to leverage inherent data correlations to generate self-supervision signals and improve data representations through pre-training methods. However, these works borrow from the Cloze task in natural language processing by masking items to have the model predict item-related information, without considering the crucial role of behaviors in recommender systems. Also, there are recent attempts to leverage the knowledge from the pretraining work on outsourced data, such as knowledge graphs (Wong et al., 2021) and language corpus (Cui et al., 2022), to enhance the recommender systems.
@@ 3. Formulation & Preliminaries
In this section, we formulate the studied problem and introduce the preliminaries. A recommendation dataset $\mathcal{D}=\{d_{z}\}_{z=1}^{N}$ is represented as a set of interactions $d_{z}$ between user $u_{z}$ and item $i_{z}$. Each interaction $d_{z}$ can be formulated as the feature-label pair $(x_{z},y_{z})$, where $x_{z}=\{x_{j}^{z}\}_{j=1}^{M}$ contains the user and item features and a label $y\in\{0,1\}$ indicating whether the user will click the item or not, i.e., click-through rate (CTR).
@@ 4. The LIFT Framework
Figure 3. The overview workflow in LIFT. In the initial phase (Stage 1), we pretrain an encoder to convert sequences into embeddings. In the subsequent phase (Stage 2), we illustrate the data flow in the figure. In Step 1, we send $x_{z}$ as a query to the retriever. In Step 2, the retriever outputs the retrieval result. In Step 3, we input the retrieval result into the encoder. In Step 4, we obtain the embeddings from the encoder. In Step 5, we send $x_{z}$ and the encoded embeddings to the predictor to obtain the final result.
@@ 4.1. Overview
There are two steps in LIFT framework: pretraining and predicting. During pretraining, an encoder $E$ is trained on a sampled $\mathcal{S}_{\text{pretrain}}$ from $\mathcal{D}_{\text{retrieval}}$ For $predicting$, the LIFT model first uses the retriever $R$ to find the top-$K$ relevant user context sequences $C_{t}$. And then, each sequence in $C_{t}$ will be encoded by $E_{\omega}$. After aggregating the information of the target user and item $x_{t}$, the target user’s history $h_{t}$, retrieved sequence $C_{t}$, the predictor $F_{\theta}$ will generate the output label prediction.
@@ 4.2. Encoder
Model Architecture. LIFT adopts a decoder-only Transformer as the encoder, which is based on the original implementation described in Vaswani et al. (2017). Transformer has been shown to perform strongly in language modeling (Brown et al., 2020; Devlin et al., 2019), machine translation (Vaswani et al., 2017), recommender system (Sun et al., 2019; Zhang et al., 2024), etc. Different tasks may leverage different parts of the Transformer architecture. The difference between the Transformer encoder and decoder is that the self-attention sub-layer in the decoder is modified to make sure that the prediction in position $p$ could only depend on the positions earlier than $p$ (Vaswani et al., 2017). For the data of the recommender system exhibiting temporal characteristics, we use the decoder part of the Transformer. We also compare the performance of different encoders in the experiment, which shows the superiority of the Transformer decoder.
@@
To feed the interaction $d_{z}$ into the encoder, $d_{z}$’s feature part $x_{z}$ is fed into an embedding layer. The embedding layer will map the one-hot encoded vector $x_{z}$ into a dense vector $\bf{x}_{z}$, where each feature $a_{i}$ in $x_{z}$ will be mapped into a $w$-dimensional vector and then be concatenated into one $(M\times w)$-dimensional dense vector. The label part of the $d_{z}$ will be preserved. In the training stage, if the interaction $d_{z}$ is chosen to be masked, the label will be replaced by a special token [MSK]. We will describe the training loss in the next subsection. After processing the sequence, the output will be sent into the decoder-only Transformer decoder sequentially, and the output is a set of hidden states $O=\{o_{1},o_{2},...,o_{T}\}$. Because only the last step in the decoder can utilize the full information in the sequence, we choose the final state $o_{T}$ as the representation of the interaction sequence $\mathbf{s}$.
Pretraining Data Preparation. The pretraining dataset $\mathcal{S}_{\text{pretrain}}$ is sampled from $\mathcal{D}_{\text{retrieval}}$. For each user $u$ in $\mathcal{U}$, we use all the interactions in $\mathcal{D}_{retrieval}$ of $u$ to assemble an interaction sequence $s_{u}$. Then we get a sequence set $S$. For any sequence $s$ in $S$, we pick up a subsequence every $L$ interactions. Using these subsequences, we get the pretraining dataset $\mathcal{S}_{\text{pretrain}}$.
Mask Behavior Loss. To leverage the inherent information contained within the context, we formulate a self-supervised loss function. In contrast to certain existing sequential recommendation models (Zhou et al., 2018b, 2019), wherein only the target interaction label is considered as the prediction target for any given interaction sequence, all other interaction labels within the sequence are omitted. It is rational to argue that incorporating more behavioral information from the context as the supervision signal would lead to a more comprehensive representation of the context. And compared to previous pretraining methods (Sun et al., 2019; Zhou et al., 2020) that mask items or item parts, we propose masking user behavior to better incorporate user behavior information into the pretrained model, enhancing user-item interaction modeling.
As shown in Figure 2, the left part describes the encoder architecture and loss in the pretraining procedure. In order to train a deep representation of the sequence, we mask parts of the interaction’s labels and then predict those masked behaviors. Unlike the Cloze task (Taylor, 1953) in natural language processing, it is difficult to predict the masked tokens in the user interaction sequences because the data in recommendation systems lacks syntactic structure like that found in natural language. An interaction $d_{i}$ in the input sequence $s$ can be represented as:
@@
We adopt the binary cross entropy loss as the objective function. In the pre-training procedure, we feed the masked interactions’ corresponding outputs as input into a multi-layer perceptron (MLP) to predict the masked labels. Thus the pretrain loss $\mathcal{L}_{\text{pretrain}}$ can be represented as $\mathcal{L}_{\text{pretrain}}=$
(10) $-\sum_{d_{i}\in\mathcal{M}}\left[y_{i}\log\sigma(\text{MLP}(o_{i}))+(1-y_{i})\log(1-\sigma(\text{MLP}(o_{i})))\right]$